OpenStudy (anonymous):

The volume (V) of a cylinder is V = π r2 h. Find h, if V = 141.3, π = 3.14, and r = 3. Round to the nearest tenth if necessary.

OpenStudy (anonymous):

You simply plug in your values here and solve for h. So: \[V=\pi r^2 h\] Plug in: \[141.3=\pi 3^3 h\] \[141.3=9 \pi h\] and solve for h.

OpenStudy (anonymous):

?